# Frostman random variables, entropy inequalities, and applications

### v4

@@ 4.2 Diagonal polynomials of higher degrees with independent random variables
$\qquad\geq n(2s+\epsilon_{1}^{\prime})P_{E}(|D_{n}(X_{1}-X_{2})|\geq\zeta_{1},|D_{n}(Y_{1}-Y_{2})|\geq\zeta_{2})$
$\qquad\qquad+ns(P_{E}(|D_{n}(X_{1}-X_{2})|<\zeta_{1}\text{ or }|D_{n}(Y_{1}-Y_{2})|<\zeta_{2}))$
$\qquad\qquad+O_{d,s}(-n(s_{1}+s_{2})+H_{n}(X_{1}-X_{2})+H_{n}(Y_{1}-Y_{2}))+O(\log K).$ (61)
Recall that $\zeta_{i}\geq 2^{-ns_{i}-4}$ $(i=1,2)$, which gives $\eta_{i}+2^{-n}\leq 20\eta_{i}$. Also note that $X_{1}-X_{2}$ and $Y_{1}-Y_{2}$ are $(s_{1},O(C_{1}))$-Frostman and $(s_{2},O(C_{2}))$-Frostman, respectively. We deduce that
$P_{E}(|D_{n}(X_{1}-X_{2})|<\zeta_{1})\leq C_{1}(20\zeta_{1})^{s_{1}}\leq\epsilon_{1}^{\prime}/16,\quad P_{E}(|D_{n}(Y_{1}-Y_{2})|<\zeta_{2})\leq C_{2}(20\zeta_{2})^{s_{2}}\leq\epsilon_{1}^{\prime}/16,$
and
$H_{n}(X_{1}-X_{2})+H_{n}(Y_{1}-Y_{2})\geq n(s_{1}+s_{2})-\log_{+}C_{1}-\log_{+}C_{2}+O(1).$ (62)
So $P_{E}(|D_{n}(X_{1}-X_{2})|\geq\zeta_{1},|D_{n}(Y_{1}-Y_{2})|\geq\zeta_{2})\geq 1-\epsilon_{1}^{\prime}/8$. Now (61) and (62) leads to
@@
$\qquad\geq n(2s+\epsilon_{1}^{\prime})(1-\epsilon_{1}^{\prime}/8)+O(\log\mathfrak{K})\geq n(2s+3\epsilon_{1}^{\prime}/4)+O(\log\mathfrak{K}).$ (63)
Finally, putting (50), (51), (4.2), (53) and (63) together, we arrive at
$H_{n}(X+Y)+7H_{n}(\rho_{1}(X)+\rho_{2}(Y))\geq n(8s+\epsilon_{1}^{\prime}/2)+O(\log\mathfrak{K}).$
By independence, we also have
$H_{n}(X+Y)\geq\frac{H_{n}(X)+H_{n}(Y)}{2}+O(1)\geq ns-\log C_{1}C_{2}+O(1).$
It follows that
$7H_{n}(X+Y)+7H_{n}(\rho_{1}(X)+\rho_{2}(Y))\geq n(14s+\epsilon_{1}^{\prime}/2)+O(\log\mathfrak{K}).$
The induction is completed by taking $\epsilon_{1}^{\prime\prime}=\epsilon_{1}^{\prime}/14$, i.e., $\epsilon_{1}^{\prime\prime}(d,s)=\epsilon_{1}^{\prime\prime}(d-1,s)/14$. ∎
@@ 4.3 Diagonal polynomials with dependent random variables
The results of Sections 4.1 and 4.2 assumed the independence of $X$ and $Y$. We now extend to the dependent case, proving Theorem 1.2 in full generality. The key tool is Theorem 3.1, which turns dependent random variables to a family of conditionally independent random variables at the cost of logarithmic error terms.
Theorem 1.2’. Let $d\geq 2$ and $\rho_{1}$, $\rho_{2}$ be real polynomials of degree $d$. Let $0<s<1$, $C,c>0$. There is a positive constant $\epsilon_{2}=\epsilon_{2}(d,s)$ such that the following statement holds: For any bivariate real random variable $(X,Y)$ that is jointly $(s_{1},s_{2},C)$-Frostman and has range contained in $c,c]^{2}$, where $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$, we have
$\begin{split}&\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\rho_{1}(X)+\rho_{2}(Y))\right\}\geq n(s+\epsilon_{2})+O(\log\mathfrak{K})\end{split}$
@@
To quantify the entropy gain beyond the trivial $ns$, we write the entropy increment as
$\Psi:=\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\rho_{1}(X)+\rho_{2}(Y))\right\}-ns.$
@@
where $c$ comes from the range of $X$ and $Y$. Let $\epsilon_{1}^{\prime\prime}=\epsilon_{1}^{\prime\prime}(d,s)$ be the constant shown in Lemma 4.8. Without loss of generality, we assume that $0<\epsilon_{1}^{\prime\prime}<1$. We use a cutoff parameter $\mathfrak{r}\geq 1$ as previous. For $i=1,2$, take $\eta_{i}=\min\Big\{\big(\frac{\epsilon_{1}^{\prime\prime}}{4\mathfrak{r}dC_{0}}\big)^{1/s_{i}},\,\frac{1}{10}\Big\}$, which satisfies $dC_{0}\eta_{i}^{s_{i}}\leq\epsilon_{1}^{\prime\prime}/4\mathfrak{r}\leq 1/4$ and $\log(1/\eta_{i})=O(\log\mathfrak{K})$.
Let $E$ be the event that $X\in\mathcal{R}(\rho_{1}^{\prime};\eta_{1})$, $Y\in\mathcal{R}(\rho_{2}^{\prime};\eta_{2})$. Since $(X,Y)$ is marginally $(s_{1},C_{0};s_{2},C_{0})$-Frostman, we deduce by Lemma 4.4 that
@@
We denote the conditional probability given $E$ by $P_{E}$. Let $(X_{0},Y_{0})$ be the random variable such that $(X_{0},Y_{0})\equiv(X,Y|E)$ (see Lemma A.2). By this construction, the functions $\rho_{1}$ and $\rho_{2}$ are weakly $(\kappa,d-1)$-bi-Lipschitz on $range_{+}(X_{0})$ and $range_{+}(Y_{0})$, respectively, with $\log\kappa=O(\log\mathfrak{K})$. ALso, the random variable $(X_{0},Y_{0})$ is both $(s_{1}+s_{2},O(C))$-Frostman, which gives
@@
where $\log K=O(\Psi)+O(n\epsilon^{\prime\prime}/\mathfrak{r})+O(\log\mathfrak{K})$.
We now apply Theorem 3.1 to $(X_{0},Y_{0})$, and obtain a quadruple $(X^{\prime\prime},Y^{\prime},X^{\prime},Y^{\prime\prime})$ with certain properties (see (i)-(iv) in Theorem 3.1). In particular, the random variables $X^{\prime\prime}$ and $Y^{\prime\prime}$ are conditionally independent given $D_{n}(X^{\prime},Y^{\prime})$. Let us take the functions $f$ and $g$ in Theorem 3.1 to be the identity functions. Then
@@
For any ball $B(z,r)$, using conditions (i)–(iv) in Theorem 3.1, we have
@@
As the expressions to follow get rather lengthy, we now introduce some shorthand. In what follows, $H_{n}(~\cdot~|x,y)$ is an abbreviation of $H(D_{n}(\cdot)~|D_{n}(X^{\prime},Y^{\prime})=(x,y)).$ By Lemma A.2, there exist random variables, depending on $(x,y)$, that have the same distributions with $(X^{\prime\prime},Y^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y))$, respectively. Similar as previously, we can apply Lemma 4.8 with these random variables, with $C_{1}=O\big(C\cdot 2^{-ns_{2}}/P_{E}(D_{n}(Y^{\prime})=y)\big)$ and $C_{2}=O\big(C\cdot 2^{-ns_{1}}/P_{E}(D_{n}(X^{\prime})=x)\big)$, to deduce that
@@
$\qquad+O_{d,s}\big(\log_{+}C-n(s_{1}+s_{2})-\log P_{E}(D_{n}(X^{\prime})=x)-\log P_{E}(D_{n}(Y^{\prime})=y)\big).$
@@
$\begin{split}&H_{n}(X^{\prime\prime}+Y^{\prime\prime}|X^{\prime},Y^{\prime})+H_{n}(\rho_{1}(X^{\prime\prime})+\rho_{2}(Y^{\prime\prime 2})|X^{\prime},Y^{\prime})\geq n(2s+\epsilon_{1}^{\prime\prime})\\ &\qquad\qquad\qquad+O_{d,s}(-n(s_{1}+s_{2})+H_{n}(X^{\prime})+H_{n}(Y^{\prime}))+O(\log\mathfrak{K}).\end{split}$ (71)
Substituting from (66), (68),(69), and further combining with $(X^{\prime},Y^{\prime})\equiv(X_{0},Y_{0})$ we obtain that
$H_{n}(X_{0})+H_{n}(Y_{0})+14\log K+O(\log\mathfrak{K})\geq n(2s+\epsilon_{1}^{\prime\prime})$ (72)
Recalling (67) and the notation $\log K=O(\Psi)+O(n\epsilon^{\prime\prime}/\mathfrak{r})+O(\log\mathfrak{K})$, the left-hand side of (72) is
$\leq 2ns+O(\Psi)+O(n\epsilon^{\prime\prime}/\mathfrak{r})+O(\log\mathfrak{K}).$ (73)
Combining (72) and (73), there is some absolute constant $\tilde{C}>0$ such that
$\tilde{C}\Psi\geq n\epsilon_{1}^{\prime\prime}+O(\log\mathfrak{K})+O(n\epsilon^{\prime\prime}/\mathfrak{r}).$
We now take $\mathfrak{r}\geq 1$ to be some absolute constant such that the last term on the right-hand side of above formula satisfies $O(n\epsilon_{1}^{\prime\prime}/\mathfrak{r})\leq n\epsilon_{1}^{\prime\prime}/2$, Therefore,
@@
for $\epsilon_{2}=\epsilon_{1}^{\prime\prime}/2\tilde{C}$. The proof is completed. ∎
@@ 4.4 More general polynomials
Now we prove Theorem 1.3 in the form of the following two theorems, which hold for more general forms.
Theorem 1.3’. Let $d\geq 2$ and $\rho_{1}$, $\rho_{2}$ be real polynomials of degree $d$. Let $0<s<1$ and $C_{1},C_{2},c>0$. There is a positive constant $\epsilon_{3}=\epsilon_{3}(d,s)$ such that the following statement holds: Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$. Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then
$\begin{split}&\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\rho_{1}(X)+\rho_{2}(Y))\right\}\\ &\qquad\qquad\qquad\geq n(s+\epsilon_{3})+O_{d,s}(\log_{+}C_{1}+\log_{+}C_{2})+O_{d}(\log_{+}c)+O_{\rho_{1},\rho_{2},s}(1).\end{split}$
for all $n\in\mathbb{N}$.
@@
Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, by Lemma 2.30 it is also jointly $(s_{1},s_{2},C_{1}C_{2})$-Frostman and marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman. Applying Theorem 1.2’, there is some $\epsilon_{2}=\epsilon_{2}(d,s)$ such that
$\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\rho_{1}(X)+\rho_{2}(Y))\right\}$
$\qquad\qquad\geq n(s+\epsilon_{2})+O_{d,s}(\log_{+}C_{1}+\log_{+}C_{2})+O_{d}(\log_{+}c)+O_{\rho_{1},\rho_{2},s}(1)$
for all $n\in\mathbb{N}$. Here the implied constant in the term $O(\log_{+}c)$ is independent of $s$, since the marginally Frostman constants do not involve $\max\{1,c\}$ as in (65). In view of $s_{1}+s_{2}\geq 2s$, Lemma 2.32(ii) shows that
$H_{n}(X,Y)\geq\frac{H_{n}(X)+H_{n}(Y)}{2}+ns-\frac{1}{2}\log C_{1}C_{2}+O(1).$
The conclusion then follows with $\epsilon_{3}=\epsilon_{2}/2$.
Theorem 1.3”. Suppose that $\phi(x,y)$ has the form (1). Let $0<s<1$ and $C_{1},C_{2},c>0$. There is a positive constants $\epsilon_{4}=\epsilon_{4}(d,\Xi,s)$ such that the following statement holds: Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$. Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then
$\begin{split}&\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\\ &\qquad\qquad\qquad\geq n(s+\epsilon_{4})+O_{d,s}(\log_{+}C_{1}+\log_{+}C_{2})+O_{d}(\log_{+}c)+O_{\phi,s}(1).\end{split}$
for all $n\in\mathbb{N}$.
Proof.
Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, Lemma 2.30 tells that it is both marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman and jointly $(s_{1},s_{2},C_{1}C_{2})$-Frostman. Then
$H_{n}(X)+H_{n}(Y)\geq H_{n}(X,Y)\geq 2ns-\log C_{1}C_{2}+O(1).$ (74)
In view of Lemma 2.32(ii), one has
$\frac{H_{n}(X)+H_{n}(Y)}{2}\leq H_{n}(X,Y)-ns+\frac{1}{2}\log C_{1}C_{2}+O(1).$
@@
Consider the random variable $(\xi_{11}X+\xi_{12}Y,\,\xi_{21}X+\xi_{22}Y)$. Its range is contained in $c^{\prime\prime},c^{\prime\prime}]^{2}$, with $c^{\prime\prime}=O_{\phi}(c)$. And Lemma 2.31, we assume without loss of generality that it is jointly $(s_{1},s_{2},C^{\prime\prime})$-Frostman, with $\log C^{\prime\prime}=\log C_{1}C_{2}+O_{\phi}(1)$, since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman. By Lemma 1.2’, there is some $\epsilon_{2}=\epsilon_{2}(d,s)>0$ such that
@@
$\begin{split}&d_{1}\cdot H_{n}(X+Y)+2d_{2}\cdot\frac{1}{2}H_{n}(X,Y)\geq H_{n}\big((\xi_{11}+\xi_{21})X+(\xi_{12}+\xi_{22})Y\big)\\ &\qquad\qquad\qquad+(d_{1}+2d_{2}-1)ns+O(\log_{+}C_{1}C_{2})+O_{\phi}(1).\end{split}$ (78)
@@
$\begin{split}&d_{1}^{\prime}\cdot H_{n}(X+Y)+2d_{2}^{\prime}\cdot\frac{1}{2}H_{n}(X,Y)\geq H_{n}(\xi_{11}X+\xi_{12}Y)\\ &\qquad\qquad\qquad+(d_{1}^{\prime}+2d_{2}^{\prime}-1)ns+O(\log_{+}C_{1}C_{2})+O_{\phi}(1).\end{split}$ (79)

### v5

@@ 4.2 Diagonal polynomials of higher degrees with independent random variables
$\qquad\geq n(2s+\epsilon_{1}^{\prime})\cdot P_{E}(|D_{n}(X_{1}-X_{2})|\geq\zeta_{1},|D_{n}(Y_{1}-Y_{2})|\geq\zeta_{2})$
$\qquad\qquad+ns\cdot P_{E}(|D_{n}(X_{1}-X_{2})|<\zeta_{1}\text{ or }|D_{n}(Y_{1}-Y_{2})|<\zeta_{2})$
$\qquad\qquad+O_{d,s}\big(-2ns+H_{n}(X_{1}-X_{2})+H_{n}(Y_{1}-Y_{2})\big)+O(\log\mathfrak{K}).$ (68)
Recall that $\zeta_{i}\geq 2^{-n-4}$ $(i=1,2)$, which gives $\zeta_{i}+2^{-n}\leq 20\zeta_{i}$. Also note that $X_{1}-X_{2}$ and $Y_{1}-Y_{2}$ are $(s_{1},O(C_{1}))$-Frostman and $(s_{2},O(C_{2}))$-Frostman, respectively. We deduce that
$P_{E}(|D_{n}(X_{1}-X_{2})|<\zeta_{1})\leq C_{1}(20\zeta_{1})^{s_{1}}\ll\epsilon_{1}^{\prime}/\mathfrak{r},\quad P_{E}(|D_{n}(Y_{1}-Y_{2})|<\zeta_{2})\leq C_{2}(20\zeta_{2})^{s_{2}}\ll\epsilon_{1}^{\prime}/\mathfrak{r}.$
So $P_{E}(|D_{n}(X_{1}-X_{2})|\geq\zeta_{1},|D_{n}(Y_{1}-Y_{2})|\geq\zeta_{2})=1+O(\epsilon_{1}^{\prime}/\mathfrak{r})$. Now (58) and (68) leads to
@@
$\qquad\geq n(2s+\epsilon_{1}^{\prime})+O_{d,s}(n\epsilon_{1}^{\prime}/\mathfrak{r}+\Psi)+O(\log\mathfrak{K}).$ (69)
Finally, putting (60) and (69) together, we arrive at
$2ns+O(n\epsilon_{1}^{\prime}/\mathfrak{r}+\Psi+\log\mathfrak{K})\geq n(2s+\epsilon_{1}^{\prime})+O_{d,s}(n\epsilon_{1}^{\prime}/\mathfrak{r}+\Psi)+O(\log\mathfrak{K}).$
That is to say, there is some constant $\mathfrak{C}=\mathfrak{C}(d,s)>0$ such that
$\mathfrak{C}\Psi\geq n\epsilon_{1}^{\prime}+O_{d,s}(n\epsilon_{1}^{\prime}/\mathfrak{r})+O(\log\mathfrak{K}).$
Now we take $\mathfrak{r}=\mathfrak{r}(d,s)\geq 1$ so that the second term on the right-hand side of above formula satisfies $O_{d,s}(n\epsilon_{1}^{\prime}/\mathfrak{r})\leq n\epsilon_{1}^{\prime}/2$. Then the induction is completed by taking $\epsilon_{1}^{\prime\prime}=\epsilon_{1}^{\prime}/2\mathfrak{C}$, i.e., $\epsilon_{1}^{\prime\prime}(d,s)=\epsilon_{1}^{\prime\prime}(d-1,s)/2\mathfrak{C}(d,s)$. ∎
@@ 4.3 Diagonal polynomials with dependent random variables
The results of Sections 4.1 and 4.2 assumed the independence of $X$ and $Y$. We now extend to the dependent case, proving Theorem 1.2 in full generality. The key tool is Theorem 3.1, which turns dependent random variables to a family of conditionally independent random variables with controllable error terms.
Theorem 1.2’. Let $d\geq 2$ and $\rho_{1}$, $\rho_{2}$ be real polynomials of degree $d$. Let $0<s<1$, $C,c>0$. There is a constant $\epsilon_{2}=\epsilon_{2}(d,s)>0$ such that the following statement holds: For any bivariate real random variable $(X,Y)$ that is jointly $(s_{1},s_{2},C)$-Frostman and has range contained in $c,c]^{2}$, where $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$, we have
$\begin{split}&\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}\big(\rho_{1}(X)+\rho_{2}(Y)\big)\right\}\geq n(s+\epsilon_{2})+O(\log\mathfrak{K})\end{split}$
@@
Similar as previous, we write
$\Psi:=\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}\big(\rho_{1}(X)+\rho_{2}(Y)\big)\right\}-ns.$
@@
where $c$ comes from the range of $X$ and $Y$. Let $\epsilon_{1}^{\prime\prime}=\epsilon_{1}^{\prime\prime}(d,s)$ be the constant shown in Lemma 4.8. Without loss of generality, we assume that $0<\epsilon_{1}^{\prime\prime}<1$. Let $\mathfrak{r}=\mathfrak{r}(d,s)\geq 1$ be a parameter to be determined later. For $i=1,2$, take $\eta_{i}=\min\Big\{\big(\frac{\epsilon_{1}^{\prime\prime}}{4\mathfrak{r}dC_{0}}\big)^{1/s_{i}},\,\frac{1}{10}\Big\}$, which satisfies $dC_{0}\eta_{i}^{s_{i}}\leq\epsilon_{1}^{\prime\prime}/4\mathfrak{r}\leq 1/4$ and $\log(1/\eta_{i})=O(\log\mathfrak{K})$.
Let $E$ be the event that $X\in\mathcal{R}(\rho_{1}^{\prime};\eta_{1})$, $Y\in\mathcal{R}(\rho_{2}^{\prime};\eta_{2})$. Since $(X,Y)$ is marginally $(s_{1},C_{0};s_{2},C_{0})$-Frostman, we deduce by Lemma 4.7(i) that
@@
We denote the conditional probability given $E$ by $P_{E}$. Let $(X_{0},Y_{0})$ be the random variable such that $(X_{0},Y_{0})\equiv(X,Y|E)$ (see Lemma A.2). By Lemma 4.7(v), the functions $\rho_{1}$ and $\rho_{2}$ are weakly $(\kappa,d)$-bi-Lipschitz on $range_{+}(X_{0})$ and $range_{+}(Y_{0})$, respectively, with $\log\kappa=O(\log\mathfrak{K})$. Also, the random variable $(X_{0},Y_{0})$ is $(s_{1}+s_{2},O(C))$-Frostman, which gives
@@
where $\log K=O(\Psi+n\epsilon_{1}^{\prime\prime}/\mathfrak{r}+\log\mathfrak{K})$.
We now apply Theorem 3.1 to $(X_{0},Y_{0})$, and obtain a quadruple $(X^{\prime\prime},Y^{\prime},X^{\prime},Y^{\prime\prime})$ such that (i) $(X^{\prime},Y^{\prime})\equiv(X_{0},Y_{0})$; (ii) $X^{\prime}$ and $X^{\prime\prime}$ are conditionally i.i.d. given $D_{n}(Y^{\prime})$; (iii) $Y^{\prime}$ and $Y^{\prime\prime}$ are conditionally i.i.d. given $D_{n}(X^{\prime})$; (iv) $(X^{\prime\prime},Y^{\prime})$ and $Y^{\prime\prime}$ are conditionally independent given $D_{n}(X^{\prime})$. In particular, the random variables $X^{\prime\prime}$ and $Y^{\prime\prime}$ are conditionally independent given $D_{n}(X^{\prime},Y^{\prime})$. Let us take the functions $f$ and $g$ in Theorem 3.1 to be the identity functions. Then
@@
For any ball $B(z,r)$, using conditions (i)–(iv) on $(X^{\prime\prime},Y^{\prime},X^{\prime},Y^{\prime\prime})$, we have
@@
As the expressions to follow get rather lengthy, we now introduce some shorthand. In what follows, $H_{n}(~\cdot~|x,y)$ is an abbreviation of $H(D_{n}(\cdot)~|D_{n}(X^{\prime},Y^{\prime})=(x,y)).$ By Lemma A.2, there exist random variables, depending on $(x,y)$, that have the same distributions with $(X^{\prime\prime},Y^{\prime\prime}|D_{n}(X^{\prime},Y^{\prime})=(x,y))$, respectively. Similar as previously, we can apply Lemma 4.8 with these random variables with $C_{1}=O\big(C\cdot 2^{-ns_{2}}/P_{E}(D_{n}(Y^{\prime})=y)\big)$ and $C_{2}=O\big(C\cdot 2^{-ns_{1}}/P_{E}(D_{n}(X^{\prime})=x)\big)$, together with the fact $s_{1}+s_{2}\geq 2s$, to deduce that
@@
$\qquad+O_{d,s}\big(\log_{+}C-2ns-\log P_{E}(D_{n}(X^{\prime})=x)-\log P_{E}(D_{n}(Y^{\prime})=y)\big).$
@@
$\begin{split}&H_{n}(X^{\prime\prime}+Y^{\prime\prime}|X^{\prime},Y^{\prime})+H_{n}(\rho_{1}(X^{\prime\prime})+\rho_{2}(Y^{\prime\prime})|X^{\prime},Y^{\prime})\geq n(2s+\epsilon_{1}^{\prime\prime})\\ &\qquad\qquad\qquad+O_{d,s}(-2ns+H_{n}(X^{\prime})+H_{n}(Y^{\prime}))+O(\log\mathfrak{K}).\end{split}$ (77)
Substituting from (74), (75), and further combining with $(X^{\prime},Y^{\prime})\equiv(X_{0},Y_{0})$, the left-hand side of (77) is bounded by
$H_{n}(X_{0})+H_{n}(Y_{0})+14\log K+O(\log\mathfrak{K}).$
Recalling (73) and the notation $\log K=O(\Psi+n\epsilon_{1}^{\prime\prime}/\mathfrak{r}+\log\mathfrak{K})$, we arrive at
$2ns+O(\Psi+n\epsilon_{1}^{\prime\prime}/\mathfrak{r}+\log\mathfrak{K})\geq n(2s+\epsilon_{1}^{\prime\prime})+O_{d,s}(\Psi+n\epsilon_{1}^{\prime\prime}/\mathfrak{r})+O(\log\mathfrak{K}).$
That is to say, there is some $\mathfrak{C}=\mathfrak{C}(d,s)>0$ such that
$\mathfrak{C}\Psi\geq n\epsilon_{1}^{\prime\prime}+O_{d,s}(n\epsilon_{1}^{\prime\prime}/\mathfrak{r})+O(\log\mathfrak{K}).$
We now take $\mathfrak{r}=\mathfrak{r}(d,s)\geq 1$ to be some constant such that the second term on the right-hand side of above formula satisfies $O_{d,s}(n\epsilon_{1}^{\prime\prime}/\mathfrak{r})\leq n\epsilon_{1}^{\prime\prime}/2$, Therefore,
@@
for $\epsilon_{2}=\epsilon_{1}^{\prime\prime}/2\mathfrak{C}$. The proof is completed. ∎
@@ 4.4 More general polynomials
We now study more general polynomials. We first prove several theorems that are weaker than Theorem 1.3, but contain explicit error terms.
Theorem 1.3’. Suppose that $\phi(x,y)$ has the form (2). Let $0<s<1$ and $C_{1},C_{2},c>0$. There is a constant $\epsilon_{3}>0$ such that the following statement holds: Let $(X,Y)$ be a bivariate real random variable whose range is contained in $c,c]^{2}$. Suppose that $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, with $1\ll_{s}s_{1},s_{2}\leq 1$ and $s_{1}+s_{2}\geq 2s$. Then
$\begin{split}&\max\left\{\frac{1}{2}H_{n}(X,Y),\,H_{n}(X+Y),\,H_{n}(\phi(X,Y))\right\}\\ &\qquad\qquad\qquad\geq n(s+\epsilon_{3})+O_{d,s}(\log_{+}C_{1}+\log_{+}C_{2})+O_{d}(\log_{+}c)+O_{\phi,s}(1)\end{split}$
for all $n\in\mathbb{N}$. Here $\epsilon_{3}=\epsilon_{3}(d,s)$ when $\phi(x,y)$ is diagonal, and $\epsilon_{3}=\epsilon_{3}(d,\Xi,s)$ when $\phi(x,y)$ is non-diagonal.
@@
Since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman, by Lemma 2.30 it is also jointly $(s_{1},s_{2},C_{1}C_{2})$-Frostman and marginally $(s_{1},C_{1};s_{2},C_{2})$-Frostman. In view of $s_{1}+s_{2}\geq 2s$, applying Lemma 2.4(iv) and Lemma 2.32 yields
$H_{n}(X)+H_{n}(Y)\geq H_{n}(X,Y)\geq 2ns-\log C_{1}C_{2}+O(1),$ (78)
and
$H_{n}(X,Y)\geq\frac{H_{n}(X)+H_{n}(Y)}{2}+ns-\frac{1}{2}\log C_{1}C_{2}+O(1).$ (79)
@@
When $\phi(x,y)$ is diagonal, we can write $\phi(x,y)=\tilde{\rho}_{1}(x)+\tilde{\rho}_{2}(y)$ with $\tilde{\rho}_{1}$, $\tilde{\rho}_{2}$ being polynomials of degree $d$. Applying Theorem 1.2’ yields $\epsilon_{2}=\epsilon_{2}(d,s)>0$ such that
$\begin{split}&\max\left\{\frac{H_{n}(X)+H_{n}(Y)}{2},\,H_{n}(X+Y),\,H_{n}(\tilde{\rho_{1}}(X)+\tilde{\rho_{2}}(Y))\right\}\\ &\qquad\qquad\geq n(s+\epsilon_{2})+O_{d,s}(\log_{+}C_{1}+\log_{+}C_{2})+O_{d}(\log_{+}c)+O_{\phi,s}(1)\end{split}$ (81)
for all $n\in\mathbb{N}$. Here the implied constant in the term $O_{d}(\log_{+}c)$ is independent of $s$, since the marginally Frostman constants do not involve $\max\{1,c\}$ as in (71). The conclusion then follows by combining (79) and (81), with $\epsilon_{3}(d,s)=\epsilon_{2}(d,s)/2$.
When $\phi(x,y)$ is non-diagonal in the form (2), we consider the random variable $(\xi_{11}X+\xi_{12}Y,\,\xi_{21}X+\xi_{22}Y)$. Its range is contained in $c^{\prime\prime},c^{\prime\prime}]^{2}$, with $c^{\prime\prime}=O_{\phi}(c)$. By Lemma 2.31, we assume without loss of generality that it is jointly $(s_{1},s_{2},C^{\prime\prime})$-Frostman, with $\log C^{\prime\prime}=\log C_{1}C_{2}+O_{\phi}(1)$, since $(X,Y)$ is conditionally $(s_{1},C_{1};s_{2},C_{2})$-Frostman. By Lemma 1.2’, there is some $\epsilon_{2}=\epsilon_{2}(d,s)>0$ such that
@@
$\begin{split}&d_{1}\cdot H_{n}(X+Y)+2d_{2}\cdot\frac{1}{2}H_{n}(X,Y)\geq H_{n}\big((\xi_{11}+\xi_{21})X+(\xi_{12}+\xi_{22})Y\big)\\ &\qquad\qquad\qquad+(d_{1}+2d_{2}-1)\big(ns-\frac{1}{2}\log_{+}C_{1}C_{2}\big)+O_{\phi}(1).\end{split}$ (84)
@@
$\begin{split}&d_{1}^{\prime}\cdot H_{n}(X+Y)+2d_{2}^{\prime}\cdot\frac{1}{2}H_{n}(X,Y)\geq H_{n}(\xi_{11}X+\xi_{12}Y)\\ &\qquad\qquad\qquad+(d_{1}^{\prime}+2d_{2}^{\prime}-1)\big(ns-\frac{1}{2}\log_{+}C_{1}C_{2}\big)+O_{\phi}(1)\end{split}$ (85)